# Seeing through Imagination: Learning Scene Geometry via Implicit Spatial World Modeling

### v1

@@ 4.2 Experimental Comparisons
Results on 3D scene understanding benchmarks. We compare our MILO against both specialized expert models and generalist MLLMs on five challenging benchmarks. As summarized in Table 1, MILO achieves superior performance across all tasks while requiring no explicit 3D point cloud encoder. Notably, compared to our direct baseline Video-3D LLM [zheng2025video], MILO shows consistent improvements across all metrics (e.g., +3.2% Acc@0.25 on ScanRefer [chen2020scanrefer], +6.2% B-4@0.5 on Scan2Cap [chen2021scan2cap]), validating the effectiveness of our implicit spatial world modeling.
@@
Results on spatial reasoning benchmarks. As shown in Table 4.1, our MILO achieves superior performance with an average accuracy of 61.7%, outperforming the baseline VG-LLM [zheng2025learning] by +2.2% absolute improvement. In addition, MILO yields consistent improvements across almost all sub-tasks, with the largest gains observed in the sub-tasks of relative direction (+5.7%), relative distance (+3.3%), and absolute distance (+3.7%). These results suggest that the proposed visual generative tuning yields more pronounced improvements on sub-tasks involving spatial direction and distance reasoning.
Results on embodied referring benchmarks. We further evaluate our MILO on the challenging RefSpatial-Bench, which contains three distinct splits of location, placement, and unseen compositional spatial relation. As reported in Table 3, MILO achieves highly competitive performance across all three splits. Compared to the RoboRefer [zhou2025roborefer] baseline, our MILO achieves absolute gains of +1.00% on the location split (48.00% vs. 47.00%) and +1.30% on the challenging unseen split.
@@ 4.3 Ablation Studies
Motivation clarification. Our approach is motivated by the visual illiteracy issue in current MLLMs, which often fail to accurately attend to relevant regions when generating responses. To systematically evaluate this, we design a grounded attention assessment on RefSpatial-Bench [zhou2025roborefer]. Specifically, we compute the attention distribution of the last response token over all visual tokens and calculate the sum of attention scores within the ground-truth mask region, termed the grounded attention score. As shown in Table 5, the proposed visual generative tuning significantly improves both the grounded attention score and the overall accuracy, demonstrating its effectiveness in enhancing the spatial awareness of MLLMs during response generation.
Ablations on visual generative tuning. We conduct ablation studies to validate the efficacy of our proposed visual generative tuning (cf. Exp #1 vs. Exp #2 in Table 4). The results consistently indicate that removing the visual generative tuning degrades performance, with a 1.7% drop in Acc@0.25 on ScanRefer [chen2020scanrefer] as a representative example.
@@
Ablations on RePE. We ablate the proposed RePE by replacing it with vanilla absolute 3D coordinate-based positional encoding [zheng2025video]. The comparative results between Exp #1 and Exp #5 in Table 4 clearly demonstrate the effectiveness of our proposed RePE, revealing a consistent performance improvement across all benchmarks.
@@ 6.1 More Experimental Results
Human evaluations of visual illiteracy issue. To validate the visual illiteracy in current MLLMs, we have conducted a grounded attention assessment on RefSpatial-Bench [zhou2025roborefer] in Section 4.3 of the main paper. Since the ScanNet series datasets [chen2020scanrefer, zhang2023multi3drefer, azuma2022scanqa, chen2021scan2cap] do not provide such affordance annotations, we design a human evaluation on the ScanQA dataset [azuma2022scanqa] to examine whether MLLMs focus on relevant visual regions when generating answers. Specifically, we randomly sample 200 instances from ScanQA and compute the attention distribution of the final response token across all visual tokens. Participants are required to judge whether the attention scores are concentrated on regions related to the question, with evaluation options including: focused, partially focused, and not focused.
As illustrated in Figure 8, our proposed MILO demonstrates a significantly higher proportion of focused attention compared to the baseline Video-3D LLM [zheng2025video]. This indicates that our proposed implicit spatial world modeling effectively mitigates visual illiteracy by directing the model’s attention to question-relevant regions.
Figure 8: Human evaluation of visual attention distribution between Video-3D LLM baseline [zheng2025video] and our MILO. Participants are required to judge whether the visual attention distributions are focused on the question-relevant regions.
@@
Figure 9: Qualitative examples of grounded attention assessment on RefSpatial-Bench [zhou2025roborefer], illustrating the attention distribution of the last response token over all the visual tokens.
@@ 6.2 More Visualization
Visualization of grounded attention assessment. In Table 5 of the main paper, we have presented the quantitative results of the grounded attention assessment on RefSpatial-Bench [zhou2025roborefer]. Here, we qualitatively visualize several examples by plotting the attention distribution of the last response token over all visual tokens. As shown in Figure 9, the baseline model RoboRefer [zhou2025roborefer] often fails to concentrate on the target regions specified in the instruction, leading to diffuse or misplaced attention. In contrast, our MILO demonstrates more focused and semantically aligned attention patterns that accurately correspond to the instructed spatial regions (e.g., the free space between objects). These visualizations provide intuitive evidence that MILO enhances the model’s grounded visual reasoning and alleviates the issue of visual illiteracy.

### v2

@@ 4.2 Experimental Comparisons
Results on 3D scene understanding benchmarks. We compare our MILO against both specialized expert models and generalist MLLMs on five challenging benchmarks. As summarized in Table 1, MILO achieves superior performance across all tasks while requiring no explicit 3D point cloud encoder. Notably, compared to our direct baseline Video-3D LLM [80], MILO shows consistent improvements across all metrics (e.g., +3.2% Acc@0.25 on ScanRefer [12], +6.2% B-4@0.5 on Scan2Cap [16]), validating the effectiveness of our implicit spatial world modeling.
@@
Results on spatial reasoning benchmarks. As shown in Table 4.1, our MILO achieves superior performance with an average accuracy of 61.7%, outperforming the baseline VG-LLM [79] by +2.2% absolute improvement. In addition, MILO yields consistent improvements across almost all sub-tasks, with the largest gains observed in the sub-tasks of relative direction (+5.7%), relative distance (+3.3%), and absolute distance (+3.7%). These results suggest that the proposed visual generative tuning yields more pronounced improvements on sub-tasks involving spatial direction and distance reasoning.
Results on embodied referring benchmarks. We further evaluate our MILO on the challenging RefSpatial-Bench, which contains three distinct splits of location, placement, and unseen compositional spatial relation. As reported in Table 3, MILO achieves highly competitive performance across all three splits. Compared to the RoboRefer [81] baseline, our MILO achieves absolute gains of +1.00% on the location split (48.00% vs. 47.00%) and +1.30% on the challenging unseen split.
@@ 4.3 Ablation Studies
Motivation clarification. Our approach is motivated by the visual illiteracy issue in current MLLMs, which often fail to accurately attend to relevant regions when generating responses. To systematically evaluate this, we design a grounded attention assessment on RefSpatial-Bench [81]. Specifically, we compute the attention distribution of the last response token over all visual tokens and calculate the sum of attention scores within the ground-truth mask region, termed the grounded attention score. As shown in Table 5, the proposed visual generative tuning significantly improves both the grounded attention score and the overall accuracy, demonstrating its effectiveness in enhancing the spatial awareness of MLLMs during response generation.
Ablations on visual generative tuning. We conduct ablation studies to validate the efficacy of our proposed visual generative tuning (cf. Exp #1 vs. Exp #2 in Table 4). The results consistently indicate that removing the visual generative tuning degrades performance, with a 1.7% drop in Acc@0.25 on ScanRefer [12] as a representative example.
@@
Ablations on RePE. We ablate the proposed RePE by replacing it with vanilla absolute 3D coordinate-based positional encoding [80]. The comparative results between Exp #1 and Exp #5 in Table 4 clearly demonstrate the effectiveness of our proposed RePE, revealing a consistent performance improvement across all benchmarks.
@@ 6.1 More Experimental Results
Human evaluations of visual illiteracy issue. To validate the visual illiteracy in current MLLMs, we have conducted a grounded attention assessment on RefSpatial-Bench [81] in Section 4.3 of the main paper. Since the ScanNet series datasets [12, 76, 2, 16] do not provide such affordance annotations, we design a human evaluation on the ScanQA dataset [2] to examine whether MLLMs focus on relevant visual regions when generating answers. Specifically, we randomly sample 200 instances from ScanQA and compute the attention distribution of the final response token across all visual tokens. Participants are required to judge whether the attention scores are concentrated on regions related to the question, with evaluation options including: focused, partially focused, and not focused.
As illustrated in Figure 8, our proposed MILO demonstrates a significantly higher proportion of focused attention compared to the baseline Video-3D LLM [80]. This indicates that our proposed implicit spatial world modeling effectively mitigates visual illiteracy by directing the model’s attention to question-relevant regions.
Figure 8: Human evaluation of visual attention distribution between Video-3D LLM baseline [80] and our MILO. Participants are required to judge whether the visual attention distributions are focused on the question-relevant regions.
@@
Figure 9: Qualitative examples of grounded attention assessment on RefSpatial-Bench [81], illustrating the attention distribution of the last response token over all the visual tokens.
@@ 6.2 More Visualization
Visualization of grounded attention assessment. In Table 5 of the main paper, we have presented the quantitative results of the grounded attention assessment on RefSpatial-Bench [81]. Here, we qualitatively visualize several examples by plotting the attention distribution of the last response token over all visual tokens. As shown in Figure 9, the baseline model RoboRefer [81] often fails to concentrate on the target regions specified in the instruction, leading to diffuse or misplaced attention. In contrast, our MILO demonstrates more focused and semantically aligned attention patterns that accurately correspond to the instructed spatial regions (e.g., the free space between objects). These visualizations provide intuitive evidence that MILO enhances the model’s grounded visual reasoning and alleviates the issue of visual illiteracy.